# The Mixed Integer Trust Region Problem

### v1

@@ Abstract
In this paper we consider the problem of minimizing a general quadratic function over the mixed integer points in an ellipsoid. This problem is strongly NP-hard, NP-hard to approximate within a constant factor, and optimal solutions can be irrational. In our main result we show that an arbitrarily good solution can be found in polynomial time, if we fix the number of integer variables. This algorithm provides a natural extension to the mixed integer setting, of the polynomial solvability of the trust region problem proven by Ye, Karmarkar, Vavasis, and Zippel. Our result removes a key bottleneck in the design and analysis of model trust region methods for mixed integer nonlinear optimization problems. The techniques introduced to prove this result are of independent interest and can be used in other mixed integer programming problems involving quadratic functions. As an example we consider the problem of minimizing a general quadratic function over the mixed integer points in a polyhedron. For this problem, we show that a solution satisfying weak bounds with respect to optimality can be computed in polynomial time, provided that the number of integer variables is fixed. It is well-known that finding a solution satisfying stronger bounds cannot be done in polynomial time, unless P=NP.
@@ 1 Introduction
Model trust region methods form an important class of algorithms for nonlinear optimization problems. The appeal of these methods lies in their general applicability, the elegant theory, and their practical success. We refer the reader to the book [4] for a thorough introduction. In model trust region methods, a quadratic function is used to “model” the nonlinear objective function, and this model is expected to be accurate (the model is “trusted”) only in a neighborhood of the current iterate $c$. To obtain the next iterate, first an affine change of variables is preformed to map $c$ to the origin, and then the model quadratic function is minimized over the neighborhood. If the 2-norm is used to define the neighborhood, the problem takes the form
@@
One key reason behind the success of model trust region methods is the fact that Equation TR can be solved efficiently. In fact, while optimal solutions of Equation TR can be irrational (see Observation 1), an arbitrarily good solution can be found in polynomial time. This celebrated result is due to Ye [24], Karmarkar [11], and Vavasis and Zippel [23], and a formal statement is given in Lemma 3. For extensions of Equation TR that can be solved in polynomial time, we direct the reader to [2] and references therein.
Model trust region methods have been introduced also for mixed integer nonlinear optimization problems (see, e.g., [9, 8, 16, 17]). For this more general type of optimization problems, if the 2-norm is used to define the neighborhood, the next iterate can be obtained by solving the mixed integer trust region (sub)problem
@@
which is the image of the set $\mathbb{Z}^{p}\times\mathbb{R}^{n-p}$ under an affine change of variables. As opposed to Equation TR, Equation MITR is strongly NP-hard and NP-hard to approximate within a constant factor, even in the pure integer case $p=n$ (see Theorem 3). Even deciding whether the feasible region of Equation MITR is nonempty, is NP-hard (see Theorem 4).
The relationship between the computational complexity of Equations TR and MITR resembles the situation in linear optimization, where linear programs can be solved in polynomial time, while mixed integer linear programs are strongly NP-hard and NP-hard to approximate within a constant factor. The key link between these two results was proven by Lenstra [13]: mixed integer linear programs can be solved in polynomial time, if the number of integer variables is fixed (but the number of continuous variables does not need to be be fixed). In our main result we provide the same type of link between Equations TR and MITR: we prove that an arbitrarily good solution for Equation MITR can be found in polynomial time, if the number $p$ of integer variables is fixed. The key difference between Lenstra’s result and ours, besides the different types of optimization problems considered, is that here we deal with approximate solutions rather than global optimal solutions, but this is unavoidable due to the fact that optimal solutions to Equations TR and MITR are generally irrational. In particular, our result removes a key bottleneck in the design and analysis of model trust region methods for mixed integer nonlinear optimization problems.
Our main result actually holds for a problem that is slightly more general than Equation MITR. We call it the ellipsoid-constrained mixed integer quadratic programming problem, and we define it as follows:
@@
Here $H$ is a symmetric matrix in $\mathbb{Q}^{n\times n}$, $Q$ is a symmetric positive definite matrix in $\mathbb{Q}^{n\times n}$, $h,c\in\mathbb{Q}^{n}$, and $p\in\{0,\dots,n\}$. There are two key differences with Equation MITR: the ball constraint has been replaced with an ellipsoid constraint, and the mixed integer lattice $\Pi_{p}(b^{1},\dots,b^{n})+\{c\}$ has been replaced with the “cleaner” mixed integer lattice $\mathbb{Z}^{p}\times\mathbb{R}^{n-p}$.
Equation MITR is essentially the special case of Equation E-MIQP where the matrix $Q$ is of the form $Q=B^{\mathsf{T}}B$ for some given invertible matrix $B\in\mathbb{Q}^{n\times n}$. In fact, under this assumption, Equation E-MIQP can be mapped to Equation MITR by performing the change of variables $y:=B(x-c)$. Viceversa, Equation MITR can always be mapped to Equation E-MIQP with the change of variables $y:=B^{-1}(x-c)$, where $B$ is the invertible matrix with columns $b^{1},\dots,b^{n}$. On the other hand, Equation MITR is indeed a special case of Equation E-MIQP, in the sense that not every Equation E-MIQP can be mapped to Equation MITR via a rational affine transformation, as shown by the $1$-dimensional ellipsoid $\left\{x\in\mathbb{R}:2x_{1}^{2}\leq 1\right\}$. This is because not all symmetric matrices $Q\in\mathbb{Q}^{n\times n}$ can be written in the form $B^{\mathsf{T}}B$ for some invertible matrix $B\in\mathbb{Q}^{n\times n}$.
To state our main results, we first define the concepts of $\epsilon$-approximate solutions and of size. Consider a general optimization problem $\inf\left\{f(x):x\in\mathcal{S}\right\}.$ Let $f_{\inf}$ and $f_{\sup}$ denote the infimum and the supremum of $f(x)$ on $\mathcal{S}$. For $\epsilon\in[0,1]$, we say that $x^{\diamond}\in\mathcal{S}$ is an $\epsilon$-approximate solution if
@@
There is an algorithm which, for any $\epsilon\in(0,1]$, finds an $\epsilon$-approximate solution to Equation E-MIQP. The running time of the algorithm is polynomial in the size of Equation E-MIQP and in $1/\epsilon$, provided that the number $p$ of integer variables is fixed.
Theorem 1 implies that it is significantly simpler to optimize a quadratic function over mixed integer points in an ellipsoid rather than over mixed integer points in a polyhedron, a behavior previously only known in the pure continuous setting [22, 18]. In turn, this suggests that it might be advantageous to design model trust region methods for mixed integer non-linear optimization problems that are based on the 2-norm, or ellipsoidal norms, rather than polyhedral norms.
Even though Equation MITR is a special case of Equation E-MIQP, it still plays a central role in our path to prove Theorem 1. In Sections 3 and 4, we study Equations TR and MITR, and prove a number of algorithmic and structural results of independent interest. The end goal of these two sections is to prove Proposition 1, which introduces an algorithm for Equation MITR that either finds two feasible solutions whose distance in objective value is close to $\max\left\{\lVert H\rVert_{2},\lVert h\rVert_{2}\right\}$, or finds a nonzero vector $d$ along which the feasible region is flat. In Section 5, we get back to the more general Equation E-MIQP. Mapping the ellipsoid defining its feasible region to a new ellipsoid that is arbitrarily close to the unit ball, allows us to employ Proposition 1 to obtain a theorem of the alternative for Equation E-MIQP. This result, presented in Proposition 2, gives an algorithm that either finds an $\epsilon$-approximate solution, or finds a nonzero vector $d$ along which the feasible region is flat. In the second case, we are able to map the intersection of the feasible region with a hyperplane, to the feasible region of a new Equation E-MIQP with one fewer integer variable, and as a result we complete our proof of Theorem 1. As we will see, this last step is where the additional level of generality of Equation E-MIQP over Equation MITR plays a key role. For more details on the ideas behind our proofs and techniques used, we refer the reader to the subsequent sections, where some proof overviews are provided to complement the complete and detailed proofs of the main results.
The techniques that we introduce to prove Theorem 1 are of independent interest and can be used in other mixed integer optimization problems involving a quadratic objective or quadratic constraints. To demonstrate this claim, we consider the mixed integer quadratic programming problem
@@
Here $H$ is a symmetric matrix in $\mathbb{Q}^{n\times n}$, $h\in\mathbb{Q}^{n}$, $W\in\mathbb{Q}^{m\times n}$, $w\in\mathbb{Q}^{m}$, and $p\in\{0,\dots,n\}$. Equation MIQP is NP-hard even in the pure continuous setting [18]. In the general mixed integer case, it is NP-hard to decide whether the feasible region of Equation MIQP is nonempty. However, this feasibility problem can be solved in polynomial time if the number $p$ of integer variables is fixed [13], thus we now focus on this case and assume $p$ is fixed in the subsequent discussion. In [5], the author gives an algorithm that finds an arbitrarily good solution to Equation MIQP in polynomial time, provided that the rank of $H$ is fixed. If the rank of $H$ is not fixed, it is not possible to find an arbitrarily good solution to Equation MIQP in polynomial time, unless P=NP [21]. In particular, we cannot hope to approximate Equation MIQP in polynomial time unless we are willing to accept an approximation factor that tends to one asymptotically as the problem gets larger [22]. In Section 7, we show how the algorithmic and structural results on Equation MITR, and in particular Proposition 1, can be used to design a polynomial time approximation algorithm for Equation MIQP that satisfies a weak approximation bound of this sort:
@@
Consider Equation MIQP and assume there exists $\bar{f}\in\mathbb{R}$ such that every feasible solution has objective value at least $\bar{f}$. There is an algorithm which finds an $\epsilon$-approximate solution to the problem, where the approximation factor is $\epsilon=1-\Theta(n^{-2})$. The running time of the algorithm is polynomial in the size of Equation MIQP, provided that the number $p$ of integer variables is fixed.
@@
We remark that the main emphasis of this paper lies on the theoretical computational complexity of the algorithms presented. We do not refer to practically efficient implementations of the algorithms, which we believe should be studied in the future. Before proceeding with the proofs of our main results, in Section 2 we discuss the hardness of the Problems TR, MITR, and E-MIQP.

### v2

@@ Abstract
In this paper we consider the problem of minimizing a general quadratic function over the mixed integer points in an ellipsoid. This problem is strongly NP-hard, NP-hard to approximate within a constant factor, and optimal solutions can be irrational. In our main result we show that an arbitrarily good solution can be found in polynomial time, if we fix the number of integer variables. This algorithm provides a natural extension to the mixed integer setting, of the polynomial solvability of the trust region problem proven by Ye, Karmarkar, Vavasis, and Zippel. As a result, our findings pave the way for designing efficient trust region methods for mixed integer nonlinear optimization problems. The techniques that we introduce are of independent interest and can be used in other mixed integer nonlinear optimization problems. As an example, we consider the problem of minimizing a general quadratic function over the mixed integer points in a polyhedron. For this problem, we show that a solution satisfying weak bounds with respect to optimality can be computed in polynomial time, provided that the number of integer variables is fixed. It is well known that finding a solution satisfying stronger bounds cannot be done in polynomial time, unless P=NP.
Key words: trust region problem; mixed integer quadratic programming; approximation algorithm; polynomial time algorithm; ellipsoid constraint
@@ 1 Introduction
Trust region methods form an important class of algorithms for nonlinear optimization problems. The appeal of these methods lies in their general applicability, the elegant theory, and their practical success. We refer the reader to the book [5] for a thorough introduction. In trust region methods, a quadratic function is used to model the nonlinear objective function, and this model is expected to be accurate (the model is “trusted”) only in a neighborhood of the current iterate $c$. To obtain the next iterate, first an affine change of variables is preformed to map $c$ to the origin, and then the model quadratic function is minimized over a neighborhood of the origin. If the 2-norm is used to define the neighborhood, the problem takes the form
@@
One key reason behind the success of trust region methods is the fact that Equation TR can be solved efficiently. In fact, while optimal solutions of Equation TR can be irrational (see Observation 3), an arbitrarily good solution can be found in polynomial time. This celebrated result is due to Ye [27], Karmarkar [14], and Vavasis and Zippel [26], and a formal statement is given in Lemma 3. Tractability results for extensions of Equation TR that include additional linear or quadratic constraints can be found, for example, in [12, 28, 3, 2, 6].
Trust region methods have been introduced also for mixed integer nonlinear optimization problems (see, e.g., [11, 10, 19, 20]). For this more general type of optimization problems, if the 2-norm is used to define the neighborhood, the next iterate can be obtained by solving the mixed integer trust region (sub)problem
@@
which is the image of the set $\mathbb{Z}^{p}\times\mathbb{R}^{n-p}$ under an affine change of variables. As opposed to Equation TR, we will see in Section 3 that Equation MITR is strongly NP-hard and NP-hard to approximate within a constant factor, even in the pure integer case $p=n$. Even deciding whether the feasible region of Equation MITR is nonempty, is NP-hard.
The relationship between the computational complexity of Equations TR and MITR resembles the situation in linear optimization, where linear programs can be solved in polynomial time, while mixed integer linear programs are strongly NP-hard and NP-hard to approximate within a constant factor. The key link between these two results was proven by Lenstra [16]: mixed integer linear programs can be solved in polynomial time, if the number of integer variables is fixed (but the number of continuous variables does not need to be fixed). In our main result we provide the same type of link between Equations TR and MITR: we prove that an arbitrarily good solution for Equation MITR can be found in polynomial time, if the number $p$ of integer variables is fixed. The key difference between Lenstra’s result and ours, besides the different types of optimization problems considered, is that here we deal with approximate solutions rather than global optimal solutions, but this is unavoidable due to the fact that optimal solutions to Equations TR and MITR are generally irrational. In particular, our findings pave the way for designing efficient trust region methods for mixed integer nonlinear optimization problems.
Ellipsoid-constrained mixed integer quadratic programming. Our main result actually holds for a problem that is more general than Equation MITR. We call it the ellipsoid-constrained mixed integer quadratic programming problem, and we define it as follows:
@@
Here $H$ is a symmetric matrix in $\mathbb{Q}^{n\times n}$, $Q$ is a symmetric positive definite matrix in $\mathbb{Q}^{n\times n}$, $h,c\in\mathbb{Q}^{n}$, and $p\in\{0,\dots,n\}$. There are two key differences with Equation MITR: the ball constraint has been replaced with a more general ellipsoid constraint, and the mixed integer lattice $\Pi_{p}(b^{1},\dots,b^{n})+\{c\}$ has been replaced with the “cleaner” mixed integer lattice $\mathbb{Z}^{p}\times\mathbb{R}^{n-p}$.
As we will see in Section 2, Equation MITR is essentially the special case of Equation E-MIQP where the matrix $Q$ is of the form $Q=B^{\mathsf{T}}B$ for some invertible matrix $B\in\mathbb{Q}^{n\times n}$. On the other hand, Equation E-MIQP is indeed more general than Equation MITR, in the sense that there are instances of Equation E-MIQP for which there exists no rational affine transformation that maps them to Equation MITR.
An approximation algorithm for Equation E-MIQP. To state our main results, we first define the concepts of $\epsilon$-approximate solutions and of size. Consider a general optimization problem $\inf\left\{f(x):x\in\mathcal{S}\right\}.$ Let $f_{\inf}$ and $f_{\sup}$ denote the infimum and the supremum of $f(x)$ on $\mathcal{S}$. For $\epsilon\in[0,1]$, we say that $x^{\diamond}\in\mathcal{S}$ is an $\epsilon$-approximate solution if
@@
There is an algorithm which, for any $\epsilon\in(0,1]$, either finds an $\epsilon$-approximate solution to Equation E-MIQP, or correctly detects that the feasible region is empty. The running time of the algorithm is polynomial in the size of Equation E-MIQP and in $1/\epsilon$, provided that the number $p$ of integer variables is fixed.
Theorem 1 implies that it is significantly simpler to optimize a quadratic function over mixed integer points in an ellipsoid rather than over mixed integer points in a polyhedron, a behavior previously only known in the pure continuous setting [25, 21]. In turn, this suggests that it might be advantageous to design trust region methods for mixed integer nonlinear optimization problems that are based on the 2-norm, or ellipsoidal norms, rather than polyhedral norms.
Even though Equation MITR is a special case of Equation E-MIQP, it still plays a central role in our path to prove Theorem 1. In Sections 4 and 5, we study Equations TR and MITR, and we prove a number of algorithmic and structural results of independent interest. The end goal of these two sections is to prove Proposition 1, which introduces an algorithm for Equation MITR that either finds two feasible solutions whose difference in objective value is at least $\max\left\{\lVert H\rVert_{2},\lVert h\rVert_{2}\right\}$, for $\epsilon$ that goes to zero, or finds a nonzero vector $d$ along which the feasible region is flat. In Section 6, we get back to the more general Equation E-MIQP. Mapping the ellipsoid defining its feasible region to a new ellipsoid that is arbitrarily close to the unit ball, allows us to employ Proposition 1 to obtain a flatness result for Equation E-MIQP. This result, presented in Proposition 2, gives an algorithm that either finds an $\epsilon$-approximate solution, or finds a nonzero vector $d$ along which the feasible region is flat. In the second case, we are able to map the intersection of the feasible region with a hyperplane, to the feasible region of a new Equation E-MIQP with one fewer integer variable, and as a result we complete our proof of Theorem 1. As we will see, this last step is where the additional level of generality of Equation E-MIQP over Equation MITR plays a key role. For more details on the ideas behind our proofs and techniques used, we refer the reader to the subsequent sections, where some proof overviews are provided to complement the complete and detailed proofs of the main results.
A weak approximation algorithm for Equation MIQP. The techniques that we introduce to prove Theorem 1 are of independent interest and can be used in other mixed integer nonlinear optimization problems. To demonstrate this claim, we consider the mixed integer quadratic programming problem
@@
Here $H$ is a symmetric matrix in $\mathbb{Q}^{n\times n}$, $h\in\mathbb{Q}^{n}$, $W\in\mathbb{Q}^{m\times n}$, $w\in\mathbb{Q}^{m}$, and $p\in\{0,\dots,n\}$. Equation MIQP is NP-hard even in the pure continuous setting [21]. In the general mixed integer case, it is NP-hard to decide whether the feasible region of Equation MIQP is nonempty. However, this feasibility problem can be solved in polynomial time if the number $p$ of integer variables is fixed [16], thus we now focus on this case and assume $p$ is fixed in the subsequent discussion. In [7], the author gives an algorithm that finds an $\epsilon$-approximate solution to Equation MIQP in time polynomial in the size of Equation MIQP and in $1/\epsilon$, provided that the rank of $H$ is fixed. In the general case, where the rank of $H$ is not fixed, no such algorithm can exist, unless P=NP [24]. In fact, unless P=NP, there is a constant $\epsilon\in(0,1)$ such that no algorithm can find an $\epsilon$-approximate solution to Equation MIQP in polynomial time [1]. As a result, we cannot hope to approximate Equation MIQP in polynomial time unless we are willing to accept an approximation factor that tends to one asymptotically as the problem gets larger [25]. In Section 8, we show how the algorithmic and structural results on Equation MITR, and in particular Proposition 1, can be used to design a polynomial time approximation algorithm for Equation MIQP that satisfies a weak approximation bound of this sort.
@@
Consider Equation MIQP and assume there exists $\bar{f}\in\mathbb{R}$ such that every feasible solution has objective value at least $\bar{f}$. There is an algorithm which either finds an $\epsilon$-approximate solution to the problem, where the approximation factor is $\epsilon=1-\Theta(n^{-2})$, or correctly detects that the feasible region is empty. The running time of the algorithm is polynomial in the size of Equation MIQP, provided that the number $p$ of integer variables is fixed.
@@
We remark that the main emphasis of this paper lies on the theoretical computational complexity of the algorithms presented. We do not refer to practically efficient implementations of the algorithms, which we believe should be studied in the future. Before proceeding with the proofs of our main results, in Section 2 we discuss the relationship between Equations MITR and E-MIQP, and in Section 3 we study the hardness of the Problems TR, MITR, and E-MIQP.
2 Relationship between Equations MITR and E-MIQP
The next result shows that Equation MITR is essentially the special case of Equation E-MIQP where the matrix $Q$ is of the form $Q=B^{\mathsf{T}}B$ for some invertible matrix $B\in\mathbb{Q}^{n\times n}$.
Observation 1.
Up to a rational affine transformation, Equation MITR is the special case of Equation E-MIQP where the matrix $Q$ is of the form $Q=B^{\mathsf{T}}B$ for some invertible matrix $B\in\mathbb{Q}^{n\times n}$.
Proof.
Consider Equation MITR and perform the change of variables $y:=B^{-1}(x-c)$, where $B$ is the invertible matrix with columns $b^{1},\dots,b^{n}$. The optimization problem that we obtain can be written in the form
$\gamma+\min$ $\quad y^{\mathsf{T}}H^{\prime}y+{h^{\prime}}^{\mathsf{T}}y$
$\operatorname{s.t.}$ $\quad(y-c^{\prime})^{\mathsf{T}}Q(y-c^{\prime})\leq 1$
$\quad y\in\mathbb{Z}^{p}\times\mathbb{R}^{n-p},$
where $\gamma:=c^{\mathsf{T}}Hc+h^{\mathsf{T}}c$, $H^{\prime}:=B^{\mathsf{T}}HB$, $h^{\prime}:=2B^{\mathsf{T}}H^{\mathsf{T}}c+B^{\mathsf{T}}h$, $Q:=B^{\mathsf{T}}B$, and $c^{\prime}:=-B^{-1}c$.
Next, consider Equation E-MIQP and assume that the matrix $Q$ is of the form $Q=B^{\mathsf{T}}B$ for some invertible matrix $B\in\mathbb{Q}^{n\times n}$. We perform the change of variables $y:=B(x-c)$ and obtain the optimization problem
$\gamma+\min$ $\quad y^{\mathsf{T}}H^{\prime}y+{h^{\prime}}^{\mathsf{T}}y$
$\operatorname{s.t.}$ $\quad y^{\mathsf{T}}y\leq 1$
$\quad y\in\Pi_{p}(b^{1},\dots,b^{n})+\{c^{\prime}\},$
where $\gamma:=c^{\mathsf{T}}Hc+{h}^{\mathsf{T}}c$, $H^{\prime}:=B^{-\mathsf{T}}HB^{-1}$, $h^{\prime}:=2B^{-\mathsf{T}}H^{\mathsf{T}}c+B^{-\mathsf{T}}h$, $b^{1},\dots,b^{n}$ are the columns of $B$, and $c^{\prime}:=-Bc$. ∎
Next, we show that Equation E-MIQP is indeed more general than Equation MITR, by presenting instances of Equation E-MIQP for which there exists no rational affine transformation that maps them to Equation MITR. Clearly, there is no rational affine transformation that maps the $1$-dimensional ellipsoid $\left\{x\in\mathbb{R}:2x_{1}^{2}\leq 1\right\}$ to the $1$-dimensional ball $\left\{x\in\mathbb{R}:x_{1}^{2}\leq 1\right\}$. However, this simple example could be written as Equation MITR, if the constraint $x^{\mathsf{T}}x\leq 1$ is replaced with $x^{\mathsf{T}}x\leq r$, for some positive $r\in\mathbb{Q}$. Next, we give a more interesting $2$-dimensional example which shows that Equation E-MIQP is also more general than the extension of Equation MITR with a constraint of the form $x^{\mathsf{T}}x\leq r$, for some positive $r\in\mathbb{Q}$.
Observation 2.
There exists no rational affine transformation that maps the $2$-dimensional ellipsoid $\left\{x\in\mathbb{R}^{2}:2x_{1}^{2}+x_{2}^{2}\leq 1\right\}$ to a $2$-dimensional ball $\left\{x\in\mathbb{R}^{2}:x_{1}^{2}+x_{2}^{2}\leq r\right\}$, for some positive $r\in\mathbb{Q}$.
Proof.
Assume, for a contradiction, that there exists a rational affine transformation as in the statement. The affine transformation is of the form $y=B(x-c)$, for some invertible matrix $B\in\mathbb{Q}^{2\times 2}$, and $c\in\mathbb{Q}^{2}$. We must have $c=0$, since the centers of the ellipsoid and of the ball are both the origin. We then have
$x^{\mathsf{T}}\begin{pmatrix}2r&0\\ 0&r\end{pmatrix}x=y^{\mathsf{T}}y=x^{\mathsf{T}}B^{\mathsf{T}}Bx.$
If we denote by $b_{ij}$ the element of $B$ in position $(i,j)$, these elements must satisfy the system
$\begin{split}&b_{11}^{2}+b_{21}^{2}=2r\\ &b_{12}^{2}+b_{22}^{2}=r\\ &b_{11}b_{12}+b_{21}b_{22}=0.\end{split}$ (1)
From the first two equations in (1), we derive
$2=\frac{b_{11}^{2}+b_{21}^{2}}{b_{12}^{2}+b_{22}^{2}}.$ (2)
From the third equation in (1), we obtain
$b_{11}^{2}b_{12}^{2}=b_{21}^{2}b_{22}^{2}.$
Since $B$ is invertible, $b_{12}$ and $b_{22}$ cannot both equal zero. If $b_{12}\neq 0$, we multiply the right hand of (2) by $b_{12}^{2}/b_{12}^{2}$ and conclude
$2=\frac{b_{11}^{2}b_{12}^{2}+b_{21}^{2}b_{12}^{2}}{b_{12}^{2}(b_{12}^{2}+b_{22}^{2})}=\frac{b_{21}^{2}b_{22}^{2}+b_{21}^{2}b_{12}^{2}}{b_{12}^{2}(b_{12}^{2}+b_{22}^{2})}=\frac{b_{21}^{2}(b_{12}^{2}+b_{22}^{2})}{b_{12}^{2}(b_{12}^{2}+b_{22}^{2})}=\frac{b_{21}^{2}}{b_{12}^{2}}.$
Hence, $\sqrt{2}$ is the rational number $\lvert b_{21}/b_{12}\rvert$, a contradiction. On the other hand, if $b_{22}\neq 0$, we multiply the right hand of (2) by $b_{22}^{2}/b_{22}^{2}$ and conclude
$2=\frac{b_{11}^{2}b_{22}^{2}+b_{21}^{2}b_{22}^{2}}{b_{22}^{2}(b_{12}^{2}+b_{22}^{2})}=\frac{b_{11}^{2}b_{22}^{2}+b_{11}^{2}b_{12}^{2}}{b_{22}^{2}(b_{12}^{2}+b_{22}^{2})}=\frac{b_{11}^{2}(b_{12}^{2}+b_{22}^{2})}{b_{22}^{2}(b_{12}^{2}+b_{22}^{2})}=\frac{b_{11}^{2}}{b_{22}^{2}}.$
Hence, $\sqrt{2}$ is the rational number $\lvert b_{11}/b_{22}\rvert$, a contradiction. ∎